Suppose that a/c = b/d where a, b, c, and d
d are real numbers. Use both arithmetic and
algebra to show that a/c = a-b/c-d is true.
1
Expert's answer
2020-05-20T19:50:06-0400

THEOREM.


If a/b=c/d then a/c=b/d




Solution.


a/b=c/d


a/b - 1=c/d - 1


a/b - b/b=c/d - d/d


(a-b) /b = (c-d)/d


(a-b) /(c-d) =b/d




In theorem b/d=a/c


(a-b) /(c-d) =a/c




The example is solved.